Equation1 a woman is 30 years older than her daughter; Equation 2 in 6 years time she will be three times her daughters age. How old is the woman using algegragic expression
Answer by Alan3354(69443)

Equation1 a woman is 30 years older than her daughter; Equation 2 in 6 years time she will be three times her daughters age. How old is the woman
-----------
W & D = ages now:
W = D + 30
W+6 = 3(D+6)
---
D+36 = 3D+18
D = 9
W = 39
==================
The easy way:
In 6 years:
30/2 = 15
--> 15 & 45